OpenStudy (anonymous):

Add or Subtract. 6/7x +1/4

OpenStudy (jim766):

you cant them together, they are not like terms

OpenStudy (anonymous):

So then is that already the answer or?

OpenStudy (jim766):

if that is all of the problem...yeah, I would just leave it. It is simplified

OpenStudy (jim766):

are you sure you wrote it down correctly?

OpenStudy (anonymous):

Yes. It is. Could you help me with 9/3x-1 - 5x/2x+3?

OpenStudy (jim766):

\[\frac{ 9 }{ 3x -1}-\frac{ 5x }{ 2x+3 }\] is that what it looks like?

OpenStudy (anonymous):

Yes.

OpenStudy (jim766):

in order to subtract fractions, we need a common denominator in this case we can use (3x-1)(2x-3) as the common denominator

OpenStudy (jim766):

in order to make this happen I have to mult the top by whatever I have to mult the bottom by to make it (3x-1)(2x+3)

OpenStudy (anonymous):

Why did you change the (2x+3) to a minus?

OpenStudy (anonymous):

Okay, now what.

OpenStudy (jim766):

sorry my computer is whacking out....no you are right it should be (2x+3)

OpenStudy (anonymous):

It's okay, but what do I do next lol?

OpenStudy (anonymous):

That was the denominator correct? So now there should be something for the numerator?

OpenStudy (jim766):

\[\frac{ 9(2x+3) }{(3x-1)(2x+3) }-\frac{ 5x(3x-1) }{ (3x-1)(2x+3) }\]

OpenStudy (jim766):

I multiplied each top by what I needed to mult the bottom by to make it our common denominator. does that make sense?

OpenStudy (anonymous):

Yes. I see what you did. like cross multiplying.

OpenStudy (jim766):

\[\frac{ 18x+27 }{ (3x-1)(2x+3) }-\frac{ 15x^2-5x }{ (3x-1)(2x+3) }\]

OpenStudy (jim766):

\[\frac{ 18x+27-(15x^2-5x) }{ (3x-1)(2x+3) }\]

OpenStudy (jim766):

\[\frac{ -15x^2+23x+27 }{ (3x-1)(2x+3) }\]

OpenStudy (anonymous):

Is that the answer or does it simplify more?

OpenStudy (zzr0ck3r):

6/7x +1/4= (4*6)/(4*7x)+(7x)/(4*7*x) = (24+28x)/(28x)

OpenStudy (zzr0ck3r):

no need for all that ^^^ this is the answer

OpenStudy (anonymous):

Okay, but I need to show my work or there's no point in me doing this. Plus, I'm not just looking for the answer. I need help understanding it.

OpenStudy (jim766):

this is a different problem

OpenStudy (zzr0ck3r):

o lol

OpenStudy (zzr0ck3r):

well this is what they want you to do to the first one

OpenStudy (jim766):

though you are correct about the 1st one

OpenStudy (zzr0ck3r):

6/7x +1/4= (4*6)/(4*7x)+(7x)/(4*7*x) = (24+28x)/(28x) factor out 6 and (6+7x)/7x

OpenStudy (anonymous):

Was the last thing you said the simplified answer to the second one?

OpenStudy (jim766):

\[\frac{ -15x^2+23x+27 }{ 6x^2+7x-3 }\]

OpenStudy (zzr0ck3r):

one the first one... we have \[\frac{6}{7x}+\frac{1}{4}\]

OpenStudy (zzr0ck3r):

our common denom is 7x*4 so multiply top and bottom of (6/7x) by 4 and top and bottom of (1/4) by 7x and you get (4*6)/(4*7x)+(7x)/(4*7*x)

OpenStudy (zzr0ck3r):

now simplify

OpenStudy (anonymous):

Okay, thank ya'll.